OpenStudy (anonymous):

Which expression completes this equation? 36 ÷ 4 – 2 • 3 = ? A.)3 • 6 – 4 ÷ 2 B.)[18 – (10 – 6)] + 1 C.)[(8 – 6) • 3] ÷ 2 D.)36 – 2 • 3

OpenStudy (anonymous):

i would say D

OpenStudy (texaschic101):

36/4 - 2 * 3 = 9 - 6 = 3 (your gonna want to find the one that also has a solution of 3 to make them equal.) Do not forget to use the order of operations or PEMDAS -- parenthesis,exponents,(multiply/divide),(add/subtract)...in that order. A) 3 * 6 - 4/2 = 18 - 2 = 16 (nope) B) [ 18 - (10 - 6)] + 1 = [18 - 4] + 1 = 14 + 1 = 15 (nope) C) [ (8 - 6) * 3] / 2 = [2 * 3] / 2 = 6/2 = 3 (BINGO) D) 36 - 2 * 3 = 36 - 6 = 30 (nope)